OpenStudy (anonymous):

((4/x-1)-(1/(x+1)))/((x/(x-1)+(1/x+1)) can someone please help me simplify this

hero (hero):

I assume you mean this: \[\large\frac{\frac{4}{x-1} - \frac{1}{x+1}}{\frac{x}{x-1}+\frac{1}{x+1}}\]

OpenStudy (anonymous):

ineeed sorry. i got it though thanks for offering though

hero (hero):

Yes, but you probably didn't do it the way that I would do it.

hero (hero):

Show me what you did.

OpenStudy (anonymous):

i multiplied the top and bottom by (x-1)(x+1) and simplified. how would you do it?

hero (hero):

Not like that...

hero (hero):

because there's no need to multiply top and bottom by anything. You can simplify it without doing that.

hero (hero):

I guess as long as you know how to do it, that's all that matters.

OpenStudy (anonymous):

(3x+5) / x^2+2x-1